OpenStudy (anonymous):

Someone wanna help a girl out? (x/5)-(5/x)=0

OpenStudy (anonymous):

hahahhah..lol @what you said before the ques!! :D

OpenStudy (turingtest):

coercion, lol what is the least common multiple of the two denominators?

OpenStudy (anonymous):

x? 5? i don't know..

OpenStudy (turingtest):

when you have two different denominators like that you can always multiply them together to get a common denominator (though if you just multiply them all it may not be the *least* common multiple) so in your case your common deonominator is \(5\times x=5x\)

OpenStudy (anonymous):

Okay then what?

OpenStudy (turingtest):

now multiply the whole equation by that

OpenStudy (turingtest):

then please show what you have left after I will help you from there

OpenStudy (anonymous):

I have 5x(x/5-5/x=0).. How do you multiply that?

OpenStudy (turingtest):

well first you can distribute on the left (I am only writing the left side right now):\[5x(\frac x5+\frac5x)=5x\cdot\frac x5+5x\cdot\frac5x\]can you simplify this?

OpenStudy (anonymous):

Dude.. You just confused me.

OpenStudy (turingtest):

do you know how to distribute?\[a(b+c)=ab+ac\]^look familiar?

OpenStudy (anonymous):

uh uh..

OpenStudy (anonymous):

how about hit and trial an x=5 and -5..case closed! ;)

OpenStudy (turingtest):

so that is what we did above a=5x b=x/5 c=5/x then we just distributed like usual

OpenStudy (turingtest):

\[5x(\frac x5+\frac5x)=5x(\frac x5)+5x(\frac5x)\]now simplify what is\[5x(\frac x5)\]?